### v3

@@ Abstract
Deep Learning, Fracture Detection, Object Detection, Medical Image Processing, Radiology, X-ray Imaging.
@@ 1 Introduction
You Only Look Once (YOLO) model [45], as one of the most important models for object detection tasks, demonstrates satisfactory performance in fracture detection [50, 19, 51]. With the introduction of YOLOv8 [24], the latest version of the YOLO models, by Ultralytics in 2023, it has been widely used in various object detection tasks. As shown in Fig. 1, GRAZPEDWRI-DX [42] is a public dataset of 20,327 pediatric wrist trauma X-ray images. This work follows the study [25] that evaluates the performance of different network models on this dataset in a fracture detection task.
Experimental results [62] show that the model performance of combining ResBlock [17] and CBAM is better than that of CBAM, and therefore we incorporate ResBlock + CBAM (ResCBAM) to the YOLOv8 architecture for experiments. After comparing the effects of different attention modules on the YOLOv8-AM model performance, we find that the GAM module has poorer gain effect on the model performance, so we propose ResGAM, which combines ResBlock and GAM, and incorporates this new attention module into the YOLOv8 architecture.
@@ 4.2 Preprocessing and Data Augmentation
In the absence of predefined training, validation, and test sets by the GRAZPEDWRI-DX dataset publisher, we perform a random division, allocating 70% to the training set, 20% to the validation set, and 10% to the test set. Specifically, the training set comprises 14,204 images (69.88%), the validation set includes 4,094 images (20.14%), and the test set comprises 2,029 images (9.98%).
@@ 4.3.1 Parameters (Params)
4.3.2 Floating Point Operations per Second (FLOPs)
@@ 4.3.4 F1 Score
Small 640 0.616 0.389 16.06 38.27 1.9
Medium 640 0.628 0.398 33.84 98.19 2.9
Large 640 0.629 0.401 53.87 196.20 4.1
Small 1024 0.632 0.390 16.06 38.27 3.0
Medium 1024 0.643 0.415 33.84 98.19 5.7
Large 1024 0.658 0.422 53.87 196.20 8.7
[…]
@@
Small 640 0.627 0.390 11.14 28.67 1.7
Medium 640 0.633 0.401 25.86 79.10 2.5
Large 640 0.640 0.415 43.64 165.44 3.9
Small 1024 0.635 0.398 11.14 28.67 2.9
Medium 1024 0.641 0.403 25.86 79.10 5.2
Large 1024 0.643 0.416 43.64 165.44 8.0
[…]
@@
Small 640 0.614 0.374 11.14 28.67 1.9
Medium 640 0.621 0.387 25.86 79.10 2.5
Large 640 0.626 0.402 43.64 165.45 3.6
Small 1024 0.621 0.387 11.14 28.67 2.7
Medium 1024 0.624 0.401 25.86 79.10 5.2
Large 1024 0.642 0.419 43.64 165.45 7.7
[…]
@@
Small 640 0.614 0.386 13.86 34.24 2.7
Medium 640 0.628 0.405 30.27 90.26 3.9
Large 640 0.640 0.412 49.29 183.53 9.4
Small 1024 0.648 0.412 13.86 34.24 4.4
Medium 1024 0.649 0.413 30.27 90.26 12.4
Large 1024 0.650 0.418 49.29 183.53 18.1
[…]
Table 3: Quantitative comparison (Precision/Recall/F1/mAP) of fracture detection on the GRAZPEDWRI-DX dataset using YOLOv8 and YOLOv8-AM models, where Inference means the total time of prediction per X-ray image, and the total time includes the preprocessing, inference, and post-processing time.
@@
4.4 Experiment Setup
We train the YOLOv8 model and different YOLOv8-AM models on the GRAZPEDWRI-DX [42] dataset. In contrast to the 300 epochs recommended by Ultralytics [24] for YOLOv8 training, the experimental results [25] indicate that the best performance is achieved within 60 to 70 epochs. Consequently, we set 100 epochs for all models training.
@@ 4.5 Experimental Results
As shown in Tables 1 and 2, the performance of the models trained using the training set with input image size of 1024 surpasses that of models trained using the training set with input image size of 640. Nevertheless, it is noteworthy that this improvement in performance is accompanied by an increase in inference time. For example, considering the ResCBAM-based YOLOv8-AM model with the large model size, the mean Average Precision at IoU 50 (mAP 50) attains 42.2% for the input image size of 1024, which is 5.24% higher than that of 40.1% obtained for the input image size of 640. However, the inference time of the model increases from 4.1ms to 8.7ms because the model size becomes larger.
Table 1 shows the model performance of the YOLOv8-AM, employing three different attention modules including ResCBAM, SA and ECA. These experimental results are obtained with different model sizes and different input image sizes. In Table 2, the model performance of the YOLOv8-AM is presented when utilizing GAM directly. Additionally, we introduce a novel approach by incorporating ResBlock and GAM, named ResGAM, to enhance the overall model performance of the YOLOv8-AM. Specifically, when the input image size is 1024 and the model size is medium, the newly introduced ResGAM demonstrates a notable enhancement in mAP for the YOLOv8-AM model based on GAM. The mAP 50 increases from 63.7% to 64.9%, providing that our proposed ResGAM positively contributes to the performance enhancement.

### v4

@@ Abstract
Deep Learning Fracture Detection Object Detection Medical Image Processing Radiology X-ray Imaging.
@@ 1 Introduction
You Only Look Once (YOLO) model [46], as one of the most important models for object detection tasks [26], demonstrates satisfactory performance in fracture detection [19, 51, 52]. With the introduction of YOLOv8 [23], the latest version of the YOLO models, by Ultralytics in 2023, it has been widely used in various object detection tasks. As shown in Fig. 1, GRAZPEDWRI-DX [43] is a public dataset of 20,327 pediatric wrist trauma X-ray images. This work follows the study [25] that evaluates the performance of different network models on this dataset in a fracture detection task.
@@
Experimental results [63] show that the model performance of combining ResBlock [17] and CBAM is better than that of CBAM, and therefore we incorporate ResBlock + CBAM (ResCBAM) to the YOLOv8 architecture for experiments. After comparing the effects of different attention modules on the YOLOv8-AM model performance, we find that the GAM module has poorer gain effect on the model performance, so we propose ResGAM, which combines ResBlock and GAM, and incorporates this attention module into the YOLOv8 architecture.
@@ 4.2 Preprocessing and Data Augmentation
In the absence of predefined training, validation, and test sets by the dataset publisher, we perform a random division, allocating 70% to the training set, 20% to the validation set, and 10% to the test set. Specifically, the training set comprises 14,204 images (69.88%), the validation set includes 4,094 images (20.14%), and the test set comprises 2,029 images (9.98%).
@@ 4.3.1 Parameters (Params)
4.3.2 Floating Point Operations (FLOPs)
@@ 4.3.4 F1 Score
4.4 Experiment Setup
We train the YOLOv8 model and different YOLOv8-AM models on the dataset [43]. In contrast to the 300 epochs recommended by Ultralytics [23] for YOLOv8 training, the experimental results [25] indicate that the best performance is achieved within 60 to 70 epochs. Consequently, we set 100 epochs for all models training.
@@ 4.4 Experiment Setup
Small 640 61.6% 38.9% 16.06 38.27 1.9
Medium 640 62.8% 39.8% 33.84 98.19 2.9
Large 640 62.9% 40.1% 53.87 196.20 4.1
Small 1024 63.2% 39.0% 16.06 38.27 3.0
Medium 1024 64.3% 41.5% 33.84 98.19 5.7
Large 1024 65.8% 42.2% 53.87 196.20 8.7
(a) ResBlock + Convolutional Block Attention Module
@@
Small 640 62.7% 39.0% 11.14 28.67 1.7
Medium 640 63.3% 40.1% 25.86 79.10 2.5
Large 640 64.0% 41.5% 43.64 165.44 3.9
Small 1024 63.5% 39.8% 11.14 28.67 2.9
Medium 1024 64.1% 40.3% 25.86 79.10 5.2
Large 1024 64.3% 41.6% 43.64 165.44 8.0
(b) Shuffle Attention
@@
Small 640 61.4% 37.4% 11.14 28.67 1.9
Medium 640 62.1% 38.7% 25.86 79.10 2.5
Large 640 62.6% 40.2% 43.64 165.45 3.6
Small 1024 62.1% 38.7% 11.14 28.67 2.7
Medium 1024 62.4% 40.1% 25.86 79.10 5.2
Large 1024 64.2% 41.9% 43.64 165.45 7.7
(c) Efficient Channel Attention
@@
Small 640 61.4% 38.6% 13.86 34.24 2.7
Medium 640 62.8% 40.5% 30.27 90.26 3.9
Large 640 64.0% 41.2% 49.29 183.53 9.4
Small 1024 64.8% 41.2% 13.86 34.24 4.4
Medium 1024 64.9% 41.3% 30.27 90.26 12.4
Large 1024 65.0% 41.8% 49.29 183.53 18.1
(b) ResBlock + Global Attention Mechanism
Table 3: Quantitative comparison (F1 Score/mAP) of fracture detection on the GRAZPEDWRI-DX dataset using YOLOv8 and YOLOv8-AM models, where Inference means the total time of prediction per X-ray image, and the total time includes the preprocessing, inference, and post-processing time.
@@ 4.5 Experimental Results
As shown in Tables 1(c) and 2(b), the performance of the models trained using the training set with input image size of 1024 surpasses that of models trained using the training set with input image size of 640. Nevertheless, it is noteworthy that this improvement in performance is accompanied by an increase in inference time. For example, considering the ResCBAM-based YOLOv8-AM model with the large model size, the mean Average Precision at IoU 50 (mAP 50) attains 42.2% for the input image size of 1024, which is 5.24% higher than that of 40.1% obtained for the input image size of 640. However, the inference time of the model increases from 4.1ms to 8.7ms because the model size becomes larger.
Table 1(c) shows the model performance of the YOLOv8-AM, employing three different attention modules including ResCBAM, SA and ECA. These experimental results are obtained with different model sizes and different input image sizes. In Table 2(b), the model performance of the YOLOv8-AM is presented when utilizing GAM directly. Additionally, we introduce a novel approach by incorporating ResBlock and GAM, named ResGAM, to enhance the overall model performance of the YOLOv8-AM. Specifically, when the input image size is 1024 and the model size is medium, the newly introduced ResGAM demonstrates a notable enhancement in mAP for the YOLOv8-AM model based on GAM. The mAP 50 increases from 63.7% to 64.9%, providing that our proposed ResGAM positively contributes to the performance enhancement.